40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 11 01:05:29 2011, kiwamu, Update, LSC, obtained Xarm PDH signal 
    Reply  Wed Jan 12 01:39:49 2011, kiwamu, Update, LSC, locked X arm 
Message ID: 4131     Entry time: Tue Jan 11 01:05:29 2011     Reply to this: 4141
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: obtained Xarm PDH signal 

[Jenne, Zach, Kiwamu]

 

 We made some efforts to lock the X arm cavity with the infrared beam.

We eventually obtained the PDH signal from a photo diode at AS port, we are still in the mid way of the lock.

The PDH signal now is going into c1lsc's ADC.

We have to make sure which digital channel corresponds to our PDH signal,

 

(what we did)

- split the LO signal, which comes from a Marconi, just before the EOM into two path.

     One is going to the EOM and the other is going to the AP table for the demodulation, The driving frequency we are using is 11MHz.

- put RF amplifiers to make the RF signal bigger. The raw signal was small, it was about ~-50 dBm in the spectrum analyzer.

   So we connected two ZLN amplifiers. Now the RF signal is at about 0 dBm

- connected the LO and RF signals to a mixer.  Additionally we put a 9.5-11.5 MHz bandpass filter at the LO path since there was some amounts of 29.5MHz due to the RF reflection at the EOM resonant box.

     After a low pass filter by SR560 the signal shows typical PDH behaviors.

- strung a BNC cable which connects the demodulated signal and c1lsc.

    In order to connect the signal to the current working ADC, we disconnected AS166_I from a whitening board and plug our cable on it.

- tried checking the digital signal but we somehow couldn't configure DAQ setting, So actually we couldn't make sure which channel corresponds to our signal.

 

ELOG V3.1.3-